OpenStudy (anonymous):

Solve using the substitution method: x^(1/2)-4x^(1/4)=-3

OpenStudy (anonymous):

Let \(\bf u=x^{\frac{1}{4}}\). After substituting this in to your given equation you get the following quadratic:\[\bf u^2-4u=-3 \implies u^2 -4u+3=0\]Solve for \(\bf u\) and then substitute that value of \(\bf u\) back in to the equation \(\bf u=x^{\frac{1}{4}}\) and solve for \(\bf x\) by taking both sides to the power of 4. @1668168

OpenStudy (anonymous):

oh...that works too

OpenStudy (anonymous):

lol ;]

OpenStudy (anonymous):

Note that you will get 2 solutions for 'x' both of which will be correct. This can be confirmed by plugging in the x-values back in to the original equation.

OpenStudy (anonymous):

Thanks both of you